OpenStudy (anonymous):

A short stop makes an error by dropping the ball. As the ball drops, its height h in feet as a function of time, t, is modeled by h(t) = -16t2 + 3. A slow motion replay of the error shows the play at half speed. What function describes the height of the ball in the replay? (Hint: The function squares the time, so half the time is also squared in the new function.)

OpenStudy (anonymous):

i got h(t) = -16(t/2)2 + 3 is it correct?

OpenStudy (anonymous):

Think so.

OpenStudy (anonymous):

h(t) = -8(t)2 + 3???

OpenStudy (anonymous):

h(t) = (-16t2 + 3)/2??

OpenStudy (anonymous):

A short stop makes an error by dropping the ball. As the ball drops, its height h in feet as a function of time, t, is modeled by h(t) = -16t2 + 3. A slow motion replay of the error shows the play at half speed. What function describes the height of the ball in the replay? (Hint: The function squares the time, so half the time is also squared in the new function.) h(t) = (-16t2 + 3)/2 h(t) = -16(t/2)2 + 3 h(t) = -8(t)2 + 3